How wide a border of uniform width should be added to a rectangle that is 8 inches by 12 inches in order to double the area?

Answer by mananth(16946)

let x be the width
Area = L*W
Area = 1288=96
double area = 192

(8+2x)(12+2x)= 192
8(12+2x)+2x(12+2x)=192
96+40x+4x^2=192
4x^2+40x-96=0
/4
x^2+10x-24=0
x^2+12x-2x-24=0
x(x+12)-2(x+12)=0
(x+12)(x-2)=0
x the width = 2 inches
m.ananth@hotmail.ca